ROTATIONAL RESOLVED SPECTRA OF TRANSITIONS INVOLVING MOTION OF METHYL GROUP OF ACETALDHEYDE IN THE FIRST ELECTRONICALLY EXCITED STATE

YUNG-CHING CHOU, CHENG-LIANG HUANG, \underline{I-CHIA CHEN}, Department of Chemistry, National Tsing Hua University, Hsinchu 30013, Taiwan; CHI-KUNG NI, and A. H. KUNG, Institute of Atomic and Molecular Sciences, Taipei,106, Taiwan.

Fluorescence excitation spectra, at resolution 0.02 cm^{−1}, in the system Ã←X̃ were recorded for acetaldehyde in a supersonic jet. We performed full rotational analysis of bands with torsional vibrational quanta up to 4. Torsional levels from near the methyl torsional barrier to beyond that barrier are assigned. Torsional sublevel A below the torsional barrier is fitted as an asymmetric rotor but the resulting value of rotational parameter A is affected significantly by the torsional motion. For the E sublevel, K doublet states split significantly with torsional quantum number v_{t}. Anomalous transitions to A sublevels are observed arising from interaction of torsion and rotation. The positions of A and E sublevels at high v_{t} state cannot be fitted with the program involving only interaction of torsion and rotation. For v_{t} = 0-2 states the A/E splitting is reversed from those in the levels with excitation in acetyl hydrogen wagging; interaction with inversion varies the splitting of torsional sublevels and the K structures. Acknowledgment: The authors gratefully acknowledge the National Science Council and Ministry of Education of Taiwan for support of this work.